Statistics South Africa reported the following data for 2012. Labor force. 17 948 000. Employment, 13 422 000, working age population, 32 786 000.
(a) Calculate unemployment?2
6. UK’s real GDP was 1 360 trillion pounds [£] in 2009 and 1 434 trillion pounds in 2010. UK’s population was 191.5 million people in 2009 and 193.3 million in 2010.
Calculate.
(a) Economic growth.3
1
Expert's answer
2020-12-08T09:23:41-0500

a. "\bold {Answer }"


Unemployment = 4,526,000 persons


"\bold {Solution}"

Unemployment

= Total labour force - Total employed

"= 17,948,000 - 13,422,000"

"\bold {= 4,526,000 \space persons}"


b. "\bold {Answer}"

Economic Growth = 4.46%


"\bold {Solution}"

Economic growth

= growth in per capita GDP


2009 per capita GDP

"= \dfrac {2009 \space GDP}{2009 \space Population}"


"=\dfrac { £1,360 \space trillion}{191.5 \space million \space people}"


"= \u00a37,101,827.67" per person



2010 per capita GDP

"= \dfrac {2010 \space GDP}{2010 \space population}"


"= \dfrac {£1,434\space trillion }{193.3 \space million \space persons}"


"= \u00a37,418,520.43"



Economic Growth

"=\dfrac {£7,418,520.43- £7,101,827.67}{£7,101,827.67}×100\%" "= \dfrac {£316,692.76}{£7,101,827.67}×100\%"


"= 4.459313\%"


"=\bold{4.46\%}"


Therefore, economic growth is 4.46%